Would like the solution of the current, divide the resistance by 100, so they are 10,8.6,5&1 (5.00 V) €2 (10.0 V) 12 R2 (8002) 14 R4 (500 (2) A A 11 A 13 A R (1000) R3 (6002)
Answer #1


E, (50) EzClou) ER2(82) Rq (5.2) A A 1 Iz 3 R(102) CA {R3 (642) hI2 = Ist. Iz I4 I4 > < I Iz I4 Using Kirchoffs law in loop For loop 2 (J, +13-1478 -107 613 = 0 =) 811 + 1413 - 4 I4 = 10 45+ 713- 4 14 = 5 - =) For loop 3 613 +514=10 Multiplying eq 0from eq 41,+9.366 - 1,5776= 5 =)I, - Ij -0.6971 AL =) I 2 = Ist Iz - I4 0,24 65A